Andrea has a particularly old bike, and she has not lubricated the chain or gears. This causes friction inside the gears as she rides. She begins coasting on flat ground at a speed of 10 m/s, and after a few seconds is only going 8 m/s. How much energy has been converted to heat?


1
Expert's answer
2022-04-19T10:01:03-0400

"E=E_1-E_2=\frac m2(v_1^2-v_2^2)=360~J."